# Mitigating Catastrophic Forgetting in Large Language Models with Self-Synthesized Rehearsal

### v1

@@ 1 Introduction
Extensive experiments on the task sequences derived from the SuperNI dataset Wang et al. (2022) demonstrate that SSR has superior or comparable performance compared to the conventional rehearsal-based approaches, with higher data utilization efficiency. Besides, experiments on AlpacaEval and MMLU Hendrycks et al. (2021) show that SSR can also effectively preserve the generalization capabilities of LLMs in general domains.
@@ 2 Related Work
Learning a sequence of datasets continually while preserving past knowledge and skills is a crucial aspect of achieving human-level intelligence. Existing approaches to continual learning can be broadly categorized into three main categories: (i) regularization-based, (ii) architecture-based, and (iii) rehearsal-based methods. Regularization techniques Kirkpatrick et al. (2017); Cha et al. (2021); Huang et al. (2021) control the extent of parameter updates during the learning process, preventing interference with previously learned tasks. Nonetheless, these methods typically rely on hyperparameters that need to be carefully tuned for optimal performance. Architecture-based approaches Xu and Zhu (2018); Huang et al. (2019); Razdaibiedina et al. (2023) often take a different approach by learning separate sets of parameters dedicated to individual tasks. This enables the model to specialize and adapt its parameters for each task, avoiding interference between tasks and preserving task-specific knowledge. However, these approaches will introduce additional training parameters, which may not be very flexible and feasible for various LLMs.
Therefore, we focus on rehearsal-based methods de Masson d'Autume et al. (2019); Rolnick et al. (2019), which are also called replay-based methods. These methods typically involve the storage of a subset of data from previous tasks. These stored data are used for future rehearsal through techniques such as experience replay Rolnick et al. (2019) and representation consolidation Bhat et al. (2022). Prior rehearsal-based approaches for language models mainly focus on using a little bit of precedent data Scialom et al. (2022); Mok et al. (2023); Zhang et al. (2023). However, these approaches often ignore discussion on real-world applications where previous data may be limited or unavailable. Although data-free knowledge distillation methods Yin et al. (2020); Smith et al. (2021) introduce auxiliary generative models for data construction, they are primarily designed for classification tasks, which may not be effective in LLMs, where a wide range of NLP tasks are involved. Additionally, training additional generative models can be challenging and time-consuming.
@@ 3 Rehearsal-Based Continual Learning
where $r$ represents the rehearsal ratio determining the percentage of sampled training instances. Finally, we use $D^{(t)}$ to fine-tune the LLM $\theta^{(t-1)}$, obtaining the updated LLM $\theta^{(t)}$. Particularly, in the first stage, we fine-tune the base LLM $\theta^{(0)}$ on $D^{(1)}=d^{(1)}$. By doing so, the catastrophic forgetting problem of LLM can be effectively alleviated, which has been verified in previous studies Scialom et al. (2022); Mok et al. (2023); Zhang et al. (2023).
@@ 4 Our Framework
In-Context Learning Based Instance Synthesis.
@@
Formally, during each training stage $t$, we first acquire $K$ demonstrations $\{(x_{k},y_{k})\}_{k=1}^{K}$. To retain previously acquired knowledge, these demonstrations can be collected from the previous instruction data $d^{(t-1)}$ or manually constructed containing similar knowledge to $d^{(t-1)}$. We concatenate all demonstrations and utilize the base LLM to generate the synthetic instance $(\hat{x},\hat{y})=\mathrm{LLM}(\mathrm{concat}_{k=1}^{K}(x_{k},y_{k});\theta^{(0)})$.
By reordering the demonstrations and sampling multiple times, we can easily obtain different synthetic instances. It should be noted that we use the base LLM $\theta^{(0)}$ rather than the latest LLM $\theta^{(t-1)}$ to conduct ICL. This is because the ICL ability of LLMs tends to exhibit a significant degradation after supervised fine-tuning (SFT) on specific tasks, as analyzed in Wang et al. (2023).
Synthetic Output Refinement.
Through the above process, we obtain a series of synthetic instances, some of which, however, may be of low quality with unreliable outputs. To address this issue, we use the latest LLM $\theta^{(t-1)}$ to refine the output of each synthetic instance: $\bar{y}=\mathrm{LLM}(\hat{x};\theta^{(t-1)})$.
By doing so, we can ensure that each refined synthetic instance $(\hat{x},\bar{y})$ retains the knowledge acquired by the latest LLM.
Rehearsal with Selected Synthetic Instances.
@@ 5.1 Setup
  • Multi-task Learning (MTL). This is the most commonly-used baseline, where all tasks are trained simultaneously.
@@
  • RandSel($r$) Scialom et al. (2022). We randomly sample $r=$1, 10}of the original instruction data for each previous task. Note that as mentioned in Scialom et al. (2022), the abilities of LLMs can be effectively preserved with $r=$ 1%.
@@ 5.3 Experiments on 10 SuperNI Tasks
To further investigate the effectiveness of SSR in longer continual learning sequences, we evaluate SSR and all baselines on 10 SuperNI tasks. Table 2 shows that SSR surpasses all rehearsal-based baselines across all metrics. Notably, SSR achieves the highest FWT value for Llama-2-7b, indicating its superior generalization capability on unseen tasks. Moreover, as illustrated in Figure 3, SSR consistently achieves better performance in terms of AR and BWT compared to rehearsal-based baselines throughout the entire continual learning process. Although SSR falls behind RandSel(10%) in terms of FWT in the early stage, it gradually strengthens its performance as the number of training stages increases, eventually surpassing RandSel(10%). Please refer to Appendix C for more details.
@@ 5.5 Analysis
Our main experiments demonstrate surprising results that rehearsal with synthetic instances may surpass those with real instances. For the comparison of real and synthetic instances, we consider the following variants: (a) RandSel (10%): Real inputs and outputs for rehearsal. (b) RandSel (10%) w/ syn. op.: Real inputs and synthetic outputs for rehearsal. Concretely, we regenerate the outputs of randomly sampled previous instances by the latest LLM, with similar operations to SSR.
Figure 6 demonstrates that RandSel (10%) with real inputs and synthetic outputs for rehearsal, outperforms RandSel (10%) utilizing only real instances. Meanwhile, SSR, which leverages both synthetic inputs and outputs for rehearsal, achieves intermediate performance, sometimes even surpassing the other two. This indicates that real instances are not always essential and appropriate for the continual learning of LLMs. As depicted in Figure 7, real instances often lead to a slower descent in loss. Therefore, they may not be conducive to optimization due to the distribution gap between distinct datasets. Conversely, synthetic instances, with lower model perplexity, embody the LLM’s real-time acquired knowledge, which aids in smoothing the data distribution and discovering better local optima for the LLM.

### v2

@@ 1 Introduction
Extensive experiments on the task sequences derived from the SuperNI dataset Wang et al. (2022) demonstrate that SSR has superior or comparable performance compared to the conventional rehearsal-based approaches, with higher data utilization efficiency. Besides, experiments on AlpacaEval and MMLU Hendrycks et al. (2021) show that SSR can also effectively preserve the generalization capabilities of LLMs in general domains. We release our code and data at https://github.com/DeepLearnXMU/SSR.
@@ 2 Related Work
Learning a sequence of datasets continually while preserving past knowledge and skills is a crucial aspect of achieving human-level intelligence. Existing approaches to continual learning can be broadly categorized into three main categories: (i) regularization-based, (ii) architecture-based, and (iii) rehearsal-based methods. Regularization techniques Kirkpatrick et al. (2017); Cha et al. (2021); Huang et al. (2021); Zhang et al. (2022) control the extent of parameter updates during the learning process, preventing interference with previously learned tasks. Nonetheless, these methods typically rely on hyperparameters that need to be carefully tuned for optimal performance. Architecture-based approaches Xu and Zhu (2018); Huang et al. (2019); Razdaibiedina et al. (2023) often take a different approach by learning separate sets of parameters dedicated to individual tasks. This enables the model to specialize and adapt its parameters for each task, avoiding interference between tasks and preserving task-specific knowledge. However, these approaches will introduce additional training parameters, which may not be very flexible and feasible for various LLMs.
Therefore, we focus on rehearsal-based methods de Masson d'Autume et al. (2019); Rolnick et al. (2019), which are also called replay-based methods. These methods typically involve the storage of a subset of data from previous tasks. These stored data are used for future rehearsal through techniques such as experience replay Rolnick et al. (2019) and representation consolidation Bhat et al. (2022). Prior rehearsal-based approaches for language models mainly focus on using a little bit of precedent data Scialom et al. (2022); Mok et al. (2023); Zhang et al. (2023b). However, these approaches often ignore discussion on real-world applications where previous data may be limited or unavailable. Although data-free knowledge distillation methods Yin et al. (2020); Smith et al. (2021) introduce auxiliary generative models for data construction, they are primarily designed for classification tasks, which may not be effective in LLMs, where a wide range of NLP tasks are involved. Additionally, similar to introducing teacher models Miao et al. (2023); Cheng et al. (2023); Huang et al. (2024), it can be challenging and time-consuming to train additional generative models. Self-distillation methods Zhang et al. (2023a) may be useful, but catastrophic forgetting of the latest LLMs and the knowledge discrepancy among LLMs from distinct stages are still inevitable challenges.
@@ 3 Rehearsal-Based Continual Learning
where $r$ represents the rehearsal ratio determining the percentage of sampled training instances. Finally, we use $D^{(t)}$ to fine-tune the LLM $\theta^{(t-1)}$, obtaining the updated LLM $\theta^{(t)}$. Particularly, in the first stage, we fine-tune the base LLM $\theta^{(0)}$ on $D^{(1)}=d^{(1)}$. By doing so, the catastrophic forgetting problem of LLM can be effectively alleviated, which has been verified in previous studies Scialom et al. (2022); Mok et al. (2023); Zhang et al. (2023b).
@@ 4 Our Framework
In-Context Learning Based Instance Synthesis
@@
Formally, during each training stage $t$, we first acquire $K$ demonstrations $\{(x_{k},y_{k})\}_{k=1}^{K}$. To retain previously acquired knowledge, these demonstrations can be collected from the previous instruction data $d^{(t-1)}$ or manually constructed containing similar knowledge to $d^{(t-1)}$. We concatenate all demonstrations and utilize the base LLM to generate the synthetic instance $(\hat{x},\hat{y})=\mathrm{LLM}(\mathrm{concat}_{k=1}^{K}(x_{k},y_{k});\theta^{(0)})$. By reordering the demonstrations and sampling multiple times, we can easily obtain different synthetic instances. It should be noted that we use the base LLM $\theta^{(0)}$ rather than the latest LLM $\theta^{(t-1)}$ to conduct ICL. This is because the ICL ability of LLMs tends to exhibit a significant degradation after supervised fine-tuning (SFT) on specific tasks, as analyzed in Wang et al. (2023).
Synthetic Output Refinement
Through the above process, we obtain a series of synthetic instances, some of which, however, may be of low quality with unreliable outputs. To address this issue, we use the latest LLM $\theta^{(t-1)}$ to refine the output of each synthetic instance: $\bar{y}=\mathrm{LLM}(\hat{x};\theta^{(t-1)})$. By doing so, we can ensure that each refined synthetic instance $(\hat{x},\bar{y})$ retains the knowledge acquired by the latest LLM.
Rehearsal with Selected Synthetic Instances
@@ 5.1 Setup
  • Multi-task Learning (MTL). This is the most commonly used baseline, where all tasks are trained simultaneously.
@@
  • RandSel($r$) Scialom et al. (2022). We randomly sample $r=$1, 10}of the original instruction data for each previous task. Note that as mentioned in Scialom et al. (2022), the abilities of language models can be effectively preserved with $r=$ 1%.
@@ 5.3 Experiments on 10 SuperNI Tasks
To further investigate the effectiveness of SSR in longer continual learning sequences, we evaluate SSR and all baselines on 10 SuperNI tasks. Table 2 shows that SSR surpasses all rehearsal-based baselines across all metrics. Moreover, as illustrated in Figure 3, SSR consistently achieves better performance in terms of AR and BWT compared to rehearsal-based baselines throughout the entire continual learning process. Although SSR with Llama-2-7b as the base LLM falls behind RandSel(10%) in terms of FWT in the early stage, it gradually strengthens its performance as the number of training stages increases, eventually surpassing RandSel(10%). Please refer to Appendix C for more details.
@@ 5.5 Analysis
Our main experiments demonstrate surprising results that rehearsal with synthetic instances may surpass those with real instances. For the comparison of real and synthetic instances, we consider the following variants: (a) RandSel (10%): Real inputs and outputs for rehearsal. (b) RandSel (10%) w/ syn. op.: Real inputs and synthetic outputs for rehearsal. Concretely, we regenerate the outputs of randomly sampled previous instances by the latest LLM, with similar operations to SSR. Figure 6 demonstrates that RandSel (10%) with real inputs and synthetic outputs for rehearsal, outperforms RandSel (10%) utilizing only real instances. Meanwhile, SSR, which leverages both synthetic inputs and outputs for rehearsal, achieves intermediate performance, sometimes even surpassing the other two. This indicates that real instances are not always essential and appropriate for the continual learning of LLMs. As depicted in Figure 7, real instances often lead to a slower descent in loss. Therefore, they may not be conducive to optimization due to the distribution gap between distinct datasets. Conversely, synthetic instances, with lower model perplexity, embody the LLM’s real-time acquired knowledge, which aids in smoothing the data distribution and discovering better local optima for the LLM.
@@
SSR vs. regularization-based and architecture-based methods
In this paper, we focus on generating synthetic rehearsal data for instruction tuning. Prior work Zhang et al. (2023b) has demonstrated that rehearsal-based approaches are generally superior to regularization-based and architecture-based ones for instruction tuning of language models. Table 5 presents experimental results using two classical regularization-based baselines (L2 and EWC) for Llama-2-7b under CL order 1, with SSR still demonstrating its superiority. Besides, these lightweight strategies can be easily combined with SSR, potentially leading to further improvements in model performance. Moreover, architecture-based approaches heavily rely on additional task-specific parameters, which may not be practical in real-world applications where the inference time of LLMs is a crucial consideration.
Model AR FWT BWT
Llama-2-7b
MTL 53.05 - -
Non-rehearsal 17.67 2.49 -44.09
[…]
L2 29.22 7.08 -28.45
EWC 15.94 2.43 -46.30
[…]
Table 5: Comparison between SSR and regularization-based methods for Llama-2-7b on 5 SuperNI tasks under CL order 1.
@@ 6 Conclusion
Acknowledgements
The project was supported by National Key R&D Program of China (No. 2022ZD0160501), National Natural Science Foundation of China (No. 62276219), and the Public Technology Service Platform Project of Xiamen (No.3502Z20231043). We also thank the reviewers for their insightful comments.